Your parents have accumulated a $140,000 nest egg. They have been planning to use this money to pay college costs to be incurred by you and your sister, Courtney. However, Courtney has decided to forgo college and start a nail salon. Your parents are giving Courtney $23,000 to help her get started, and they have decided to take year-end vacations costing 59,000 per year for the next four years. Use 9 percent as the appropriate interest rate throughout this problem. Use Appendx A and Appendix D for an approximate answer, but calculate your f answer final using the formula and financial calculator methods. a. How much money will your parents have at the end of four years to help you with graduate school, which you will start then? (Round your final answer to 2 decimal places.) Funds available for graduate school b. You plan to work on a master's and perhaps a PhD. Igraduate school costs $25,100 per year, approximately how long will you be able to stay in school based on these funds? (Round your final answer to 2 decimal places.) Number of years
10 17 45 17 ''2334579-14 18 22 28 35 44 55 $ 800 200 Ano 000 ser 500 013 000 34 64 aru e 7 140 004 hca 200 400 350 58 13 31 74 18 22 26 31 38 95 201 005 500 400 500 112 001 250 79 2 2 125 201 1-105 2 2 2. 750 2 199 360 ENT 012 -rza-202 200-004 1 000 000 2 2 2 2 250 200 004 006 997 202 200 120 220 001 011 198 ANI 27 007 10 12 200 140 400 180 77 001 4 ese 72 7 10 12 14 16 40 170 170 198 one 197 200 201 200 013 202 86 7 789 10 12 13 25 50 aa 125 140 170 195 210 201 200 ZN 210 ara 400腦矧崩201 006 Akt 001 002 012 ies 004 105 110 125 198 49 101 194 96 008 220 202 123456789 10 1 2 3 4 5 6 7 8 9 20 25 30 40 50 12", 4 5 6 7 8 g 10 1 2 3 4 15 16 17 18 19 20 25 30 40 50
3026.808 22.306 19.600 17.292 15.372 13.765 12.40 11258 10274 9427 8694 8.055 4032.836 27.35 23115 19.793 17.159 15.046 3.33 11.926 10757 9.779 89618244 003.190 31.424 25730 21.482 18250 15.702 13.80 12.233 10902 9.91 90 304 1,668 17 1605 .585 586 1.5471526 1381 1224 .111 2.361 2322 2.2R3 2246 2.210 2174 2.140 2106 1.952 1816 1A96 15A9 1.407 4 2.974 2914 2.B56 2.798 2.743 2600 2.600 2589 2.362 2.16ยิ 1.007 1849 1.606 5. 3.517 3.433 a352 3.274 3.199 3.127 3.058 2.991 2.689 2.486 2.220 2.035 1.737 9 0.132 4.945 4.772 4.607 4.451 4.303 4.163 4.001 3.463 3.019 2.000 2.379 1.948 10 5.426 5216 5.019 4833 4.ย59 4494 4.309 4.182 3.571 3.092 2.715 2.414 1.965 5.887 5453 5.234 5.02 438 856 4484327 3.5 3147 2.752 2438 1.977 5.918 5660 5.421 197 4.98ล 4793 4.611 4.439 3.725 3190 2.779 ネ456 1.986 13 6.122 5842 5,553 5342 5.118 4.010 4.715 4.533 3.780 3.223 2.79ฉ 2.489 1.000 16 6.604 265 5.064 5.668 5.406 162 4.008 4.730 3.887 3.283 2834 2.489 1.007 18.840 .467 .128 5818 .534 523 .033 4812 3.928 3.304 2.8442494 1.9 025 8623 .25 5929 .620 5353 5.10 4870 3.54 3316 2.850 247 .9 7330 6873 6.46 6007 5.768 5467 5.196 4948 3 3329 2668 2 2000 307406 7003 6.566 6177 9 517235 497 30 3332 2.867 2500 2.000 407.634 7.105 6.642 6233 5.871 58 6.258 4997 3.9 3333 2.857 2500 2.000 50776 7.133 6.061 24 5.880 55 6.22 499 4.000 3333 2.867 2.500 2.000
Answer #1

a)

amount available at beginning = 140000-23000 = 117000

amount available after year 1 end = 117000 * 1.09 = 127530

amount available after year 2 beginning = 127530 -9000 = 118,530

amount available after year 2 end = 118,530 * 1.09 = 129,197.70

amount available after year 3 beginning = 129,197.70 -9000 = 129,197.70

amount available after year 3 end = 129,197.70* 1.09 = 131,015.49

amount available after year 4 beginning = 131,015.49 -9000 =122,015.49

amount available after year 4 end = 122,015.49* 1.09 = 132,996.88

amount available after year 5 beginning = 132,996.88 -9000 = 123,996.88

money parents have at the end of four years = 123,996.88...........ans

-------
b)

If the school fees is $25,100 per year then one can study $123,996.88/$25,100 years i.e 4.94 year in that school.
Answer #2

a of the above is correct but b is incorrect
b) should be continued the same as A.
money parents have at the end of 4 years 123996.88 X 1.09 - 25100
continue repeating calculations until there is a remainder less than school fees and divide it by 25100 and add it to the # of times you subtracted the school fees

source: intermediate accounting
answered by: daniel